# Discrete-to-continuum limits of semilinear stochastic evolution equations in Banach spaces

### v1

@@ 6.3. Global well-posedness and convergence for dissipative drifts
In this section, we consider a class of equations whose drift coefficients satisfy not only (F1-B) (which would only guarantee local well-posedness), but also a type of dissipativity condition, also used in [56], which allows us to deduce global existence.
@@
in order to establish that (F1′′-B) holds for $b^{\prime}=0$. The next example shows how (6.20) can be verified in our situation of main interest. It is an elaborated version of [56, Example 4.2].
@@
Since the inequality is preserved under convex combinations and weak limits of $x_{n}^{*}$, the above characterization of $\partial\|u_{n}\|_{B_{n}}$ shows that we only need to check it for $x_{n}^{*}=r\delta_{\hat{\xi}}$, where $r\in\operatorname{sgn}u_{n}(\hat{\xi})$ for $\hat{\xi}\in\mathcal{M}_{n}$ such that $\|u_{n}\|_{B_{n}}=\lvert u_{n}(\hat{\xi})\rvert$. That is, it suffices that
@@ 7. Outlook
where $\Gamma$ is a convex subset of a (Hilbertian) state space $H$ and the functional $I_{\Gamma}\colon H\to(-\infty,\infty]$ vanishes on $\Gamma$ and equals $\infty$ outside of it. In the first work on this problem, Nualart and Pardoux [66] used a direct approach to show existence and uniqueness in a setting which corresponds to $H\coloneqq L^{2}(0,1)$, $A\coloneqq-\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}x^{2}}$ with homogeneous Dirichlet or Neumann boundary conditions, and $\Gamma\coloneqq K_{0}$, where
@@
The result for higher dimensions $m\in\mathbb{N}$ can be derived from the $m=1$ case. Indeed, by separation of variables in the continuum case, or by writing the discretized operator as a Kronecker sum of $m$ one-dimensional discretizations, one finds that the eigenvalues and eigenvectors of the $m$-dimensional operators are sums and products, respectively, of their $1$-dimensional counterparts. From this, one can deduce the desired result. ∎
@@
a Replace (A.12) by the identity $\|U\|_{\mathscr{L}_{2}(L^{2}(\mathcal{M}))}^{2}=\sum\nolimits_{j}\lvert\alpha_{j}\rvert^{2}\|f_{j}\|_{L^{2}(\mathcal{M})}^{2},$ which follows directly from the definition of the the Hilbert–Schmidt norm. Assumption 3.8 is not required since all the eigenfunctions are $L^{2}$-normalized. The sufficiency of Assumption 3.5(i) and the $\widetilde{\mathbb{P}}$-a.s. convergence in the conclusion are due to the use of Theorem 3.12a instead of Theorem 3.12b.
@@
b The differences with part a are the use of Theorem 3.13c instead of 3.13b , and the fact that Assumption 3.8 is automatically satisfied. ∎
@@ Appendix B Fractional parabolic integration
For a sequence of operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on Banach spaces which satisfy some of the discrete-to-continuum assumptions from the main text, the proposition above implies the following corollary regarding uniform boundedness of the sequence (~$(\widetilde{\mathfrak{I}}_{A_{n}}^{s})_{n\in\overline{\mathbb{N}}}$, where ~$\widetilde{\mathfrak{I}}_{A_{n}}^{s}\coloneqq\Lambda_{n}\mathfrak{I}_{A_{n}}^{s}\Pi_{n}$ for all $n\in\overline{\mathbb{N}}$. From this, in turn, one can derive Proposition B.3 below asserting the strong convergence of these operators.
@@ Appendix C Uniformly sectorial sequences of operators
We first recall the concept of sectorial operators. Given $\omega\in(0,\pi)$, we say that a linear operator $A\colon\mathsf{D}(A)\subseteq E\to E$ on a (real or complex) Banach space $E$, with spectrum $\sigma(A):=\mathbb{C}\setminus\rho(A)$, is said to be $\omega$-sectorial if
@@
Let the operators $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ on the Banach spaces $(E_{n},\|\,\cdot\,\|_{E_{n}})_{n\in\overline{\mathbb{N}}}$ be uniformly sectorial of angle $\omega\in[0,\frac{1}{2}\pi)$, and denote by $(S_{n}(t))_{t\geq 0}$ the bounded analytic $C_{0}$-semigroups generated by $-A_{n}$. Let Assumptions 2.1 and 2.2 be satisfied with $w=0$, and suppose that the equivalent statements a and b in Theorem 2.3 hold. Then we have, for all $\alpha\in(0,\infty)$, $x\in E_{\infty}$ and $0<a<b<\infty$,
@@
$A_{n}^{\alpha}S_{n}(t)=\frac{1}{2\pi i}\int_{\partial\Sigma_{\omega^{\prime}}}z^{\alpha}e^{-tz}R(z,A_{n})\mathop{}\!\mathrm{d}z$ (C.6)
To see this, define the functions $f_{\alpha},g_{t}\colon\Sigma_{\omega^{\prime}}\to\mathbb{C}$ by $f_{\alpha}(z)\coloneqq z^{\alpha}$ and $g_{t}(z)\coloneqq e^{-tz}$ for $z\in\Sigma_{\omega^{\prime}}$ and $t\in(0,\infty)$. Denote by $f_{\alpha}(A_{n})$ and $g_{t}(A_{n})$ the operators obtained via the extended Dunford calculus for sectorial operators as defined in [50, Definition 15.1.8]. Then $f_{\alpha}(A_{n})$ is the fractional power $A_{n}^{\alpha}$ in the sense of [50, Definition 15.2.2], which satisfies $f_{\alpha}(A_{n})x=f_{\alpha}(\lambda)x=\lambda^{\alpha}x$ if $A_{n}x=\lambda x$, hence this (more general) definition agrees with the spectrally defined fractional powers in the setting of (3.13). Moreover, $g_{t}(A)=S(t)$ by [50, Theorem 15.1.7]. Since $S(t)$ is bounded, we have $(f_{\alpha}g_{t})(A_{n})=f_{\alpha}(A_{n})g_{t}(A_{n})$ by [50, Proposition 15.1.12]. Finally, the function $(f_{\alpha}g_{t})(z)=z^{\alpha}e^{-z}$ is holomorphic and has (super)polynomial decay at 00 and $\infty$, and thus belongs to the domain of the primary Dunford calculus [50, Definition 15.1.1], so that $(f_{\alpha}g_{t})(A_{n})=\frac{1}{2\pi i}\int_{\partial\Sigma_{\omega^{\prime}}}z^{\alpha}e^{-tz}R(z,A_{n})\mathop{}\!\mathrm{d}z$. Putting all these observations together yields (C.6). Applying the projection and lifting operators and parametrizing the complex integral yields, for all $x\in E_{\infty}$,

### v2

@@ 6.3. Global well-posedness and convergence for dissipative drifts
In this section, we consider a class of equations whose drift coefficients satisfy not only (F1-B) (which would only guarantee local well-posedness), but additionally a type of dissipativity condition, also used in [56], which allows us to deduce global existence.
@@
in order to establish that (F1′′-B) holds for $b^{\prime}=0$. The next example shows how the relation (6.20) can be verified in our situation of main interest. It is an elaborated version of [56, Example 4.2].
@@
Since the inequality is preserved under convex combinations and weak limits of $x_{n}^{*}$, the above characterization of $\partial\|u_{n}\|_{B_{n}}$ shows that it only needs to be checked for $x_{n}^{*}=r\delta_{\hat{\xi}}$, where $r\in\operatorname{sgn}u_{n}(\hat{\xi})$ for $\hat{\xi}\in\mathcal{M}_{n}$ such that $\|u_{n}\|_{B_{n}}=\lvert u_{n}(\hat{\xi})\rvert$. That is, it suffices to verify that
@@ 7. Outlook
where $\Gamma$ is a convex subset of a (Hilbertian) state space $H$ and the indicator functional $I_{\Gamma}\colon H\to(-\infty,\infty]$ is defined to vanish on $\Gamma$ and equal $\infty$ outside of it. In the first work on this problem, Nualart and Pardoux [66] used a direct approach to show existence and uniqueness in a setting which corresponds to $H\coloneqq L^{2}(0,1)$, $A\coloneqq-\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}x^{2}}$ with homogeneous Dirichlet or Neumann boundary conditions, and $\Gamma\coloneqq K_{0}$, where
@@
Conflict of interest and data availability statement
The authors declare that they have no conflict of interest. Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
@@
The result for higher dimensions $m\in\mathbb{N}$ can be derived from the $m=1$ case. Indeed, by separation of variables in the continuum case, or by writing the discretized operator as a Kronecker sum of $m$ one-dimensional discretizations, one finds that the eigenvalues and eigenvectors of the $m$-dimensional operators are given by sums and products, respectively, of their $1$-dimensional counterparts. From this, one can deduce the desired result. ∎
@@
a Replace (A.12) by the identity $\|U\|_{\mathscr{L}_{2}(L^{2}(\mathcal{M}))}^{2}=\sum\nolimits_{j}\lvert\alpha_{j}\rvert^{2}\|f_{j}\|_{L^{2}(\mathcal{M})}^{2},$ which follows directly from the definition of the Hilbert–Schmidt norm. Assumption 3.9 is not required since all the eigenfunctions are $L^{2}$-normalized. The sufficiency of Assumption 3.5(i) and the $\widetilde{\mathbb{P}}$-a.s. convergence in the conclusion are due to the use of Theorem 3.13a instead of Theorem 3.13b.
@@
b The differences with part a are the use of Theorem 3.14c instead of 3.14b, and the fact that Assumption 3.9 is automatically satisfied. ∎
@@ Appendix B Fractional parabolic integration
For a sequence of operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on Banach spaces which satisfy the appropriate discrete-to-continuum assumptions from the main text, the proposition above implies the following corollary regarding uniform boundedness of the sequence (~$(\widetilde{\mathfrak{I}}_{A_{n}}^{s})_{n\in\overline{\mathbb{N}}}$, where ~$\widetilde{\mathfrak{I}}_{A_{n}}^{s}\coloneqq\Lambda_{n}\mathfrak{I}_{A_{n}}^{s}\Pi_{n}$ for all $n\in\overline{\mathbb{N}}$. From this, in turn, one can derive Proposition B.3 below asserting the strong convergence of these operators.
@@ Appendix C Uniformly sectorial sequences of operators
We first recall the concept of sectorial operators. Given $\omega\in(0,\pi)$, we say that a linear operator $A\colon\mathsf{D}(A)\subseteq E\to E$ on a (real or complex) Banach space $E$, with spectrum $\sigma(A)\coloneqq\mathbb{C}\setminus\rho(A)$, is said to be $\omega$-sectorial if
@@
Let the linear operators $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ on the Banach spaces $(E_{n},\|\,\cdot\,\|_{E_{n}})_{n\in\overline{\mathbb{N}}}$ be uniformly sectorial of angle $\omega\in[0,\frac{1}{2}\pi)$, and denote by $(S_{n}(t))_{t\geq 0}$ the bounded analytic $C_{0}$-semigroups generated by $-A_{n}$. Let Assumptions 2.1 and 2.2 be satisfied with $w=0$, and suppose that the equivalent statements a and b in Theorem 2.3 hold. Then we have, for all $\alpha\in(0,\infty)$, $x\in E_{\infty}$ and $0<a<b<\infty$,
@@
$A_{n}^{\alpha}S_{n}(t)=\frac{1}{2\pi i}\int_{\partial\Sigma_{\omega^{\prime}}}z^{\alpha}e^{-tz}R(z,A_{n})\mathop{}\!\mathrm{d}z.$ (C.6)
To see this, define the functions $f_{\alpha},g_{t}\colon\Sigma_{\omega^{\prime}}\to\mathbb{C}$ by $f_{\alpha}(z)\coloneqq z^{\alpha}$ and $g_{t}(z)\coloneqq e^{-tz}$ for $z\in\Sigma_{\omega^{\prime}}$ and $t\in(0,\infty)$. Denote by $f_{\alpha}(A_{n})$ and $g_{t}(A_{n})$ the operators obtained via the extended Dunford calculus for sectorial operators as defined in [50, Definition 15.1.8]. Then $f_{\alpha}(A_{n})$ is the fractional power $A_{n}^{\alpha}$ in the sense of [50, Definition 15.2.2], which satisfies $f_{\alpha}(A_{n})x=f_{\alpha}(\lambda)x=\lambda^{\alpha}x$ if $A_{n}x=\lambda x$, hence this (more general) definition agrees with the spectrally defined fractional powers in the setting of (3.13). Moreover, $g_{t}(A_{n})=S_{n}(t)$ by [50, Theorem 15.1.7]. Since $S_{n}(t)$ is bounded, we have $(f_{\alpha}g_{t})(A_{n})=f_{\alpha}(A_{n})g_{t}(A_{n})$ by [50, Proposition 15.1.12]. Finally, the function $(f_{\alpha}g_{t})(z)=z^{\alpha}e^{-z}$ is holomorphic and has (super)polynomial decay at 00 and $\infty$, and thus belongs to the domain of the primary Dunford calculus [50, Definition 15.1.1], so that $(f_{\alpha}g_{t})(A_{n})=\frac{1}{2\pi i}\int_{\partial\Sigma_{\omega^{\prime}}}z^{\alpha}e^{-tz}R(z,A_{n})\mathop{}\!\mathrm{d}z$. Putting all these observations together yields (C.6). Applying the projection and lifting operators and parametrizing the complex integral yields, for all $x\in E_{\infty}$,